A water sample contains following impurities in ppm

CaSO4 = 12 ppm, Ca(HCO3)2= 10 ppm, Mg(HCO3)2 = 8.5 ppm, MgSO4 = 14 ppmCalculate temporary hardness, permanent hardness and total hardness of water sample


1
Expert's answer
2021-06-29T00:39:16-0400

Temporary hardness [Mg (HCO3)2 + Ca (HCO3)2] = 8.5 + 10 = 18.5 ppm

Permanent hardness [MgSO4+ CaSO4] = 14 + 12 = 26 ppm

Total hardness = Temporary hardness+ Permanent hardnes = 18.5 + 26 = 44.5 ppm